Below are the average heights for American boys in 1990. Age (years) Height (cm) birth 50.8 2 83.8 3 91.4 5 106.6 7 119.3 10 137.1 14 157.5 a.) Calculate the least squares line. Put the equation in the form of: ŷ = a + bx. (Round your answers to three decimal places.) B.) Find the estimated average height for a one-year-old. (Use your equation from part (d). Round your answer to one decimal place.) C.) Find the estimated average height for a eleven-year-old. (Use your equation from part (d). Round your answer to two decimal places.) D.) Use the least squares line to estimate the average height for a fifty-four-year-old man. (Use your equation from part (d). Round your answer to one decimal place.) E.) What is the slope of the least squares (best-fit) line? (Round your answer to three decimal places.) F.) nterpret the slope. (Round your answer to three decimal places.) As age increases by one year, the average height increases by how many centimeters?
Answer #1

a) Height = 63.111 + 7.255 * Age

B) If age = 1 years old then

Height = 63.111+ 7.255 * 1 = 70.4 cm

C) If age = 11 years old then

Height = 63.111 + 7.255 * 11 = 142.92 cm

D) If age = 54 years old then

Height = 63.111 + 7.255 * 54 =454.9 cm ( which impossible)

E) Slope = 7.255

F) As the age is increases by 1 year then average height increases by 7.255 cm. PL??